# Convergence of patterned matrices with random walk entries

### v1

@@ 3.2 Joint convergence
$A_{n}(t)-A_{n}(s)\coloneqq\big(S_{i,j}^{(n)}(t)-S_{i,j}^{(n)}(s)\big)_{1\leq i,j\leq n},\ n\geq 1.$
As $\{S_{i,j}^{(n)}(t)\}_{t\geq 0}$’s have independent increments, the process $\{A_{n}(t)\}_{t\geq 0}$ also has independent increments.
@@
(ii) If $L(i,j)=(\max\{i,j\},\min\{i,j\})$ then (3.7) holds and the collection $\{A(t)\}_{t\geq 0}$ in $\tilde{\mathcal{A}}$ is a FBM. Similarly, if L(i,j)=n/2|n/2|ij||L(i,j)=n/2-|n/2-|i-j||, $1\leq i,j\leq n$ then the joint moments of $\{A(t)\}_{t\geq 0}$ agree with those of a SBM.
@@
$S_{\pi}^{(n)}(\mathbb{T}_{p})=\mathbb{E}(S_{L(\pi(0),\pi(1))}^{(n)}(t_{r_{1}})-S_{L(\pi(0),\pi(1))}^{(n)}(t_{r_{1}-1}))\dots(S_{L(\pi(l-1),\pi(l))}^{(n)}(t_{r_{l}})-S_{L(\pi(l-1),\pi(l))}^{(n)}(t_{r_{l}-1})).$
@@
Note that if any index $r_{i}$, $i=1,\dots,l$ appears only once or circuit $\pi$ is non matched then by independent increments property and zero mean of $\{S_{i,j}^{(n)}(t)\}_{t\geq 0}$, we get $S_{\pi}^{(n)}(\mathbb{T}_{p})=0$. Thus, $S_{\pi}^{(n)}(\mathbb{T}_{p})=0$ for all non indexed matched circuits $\pi$. Let us write $M_{L(i,j)}^{(n)}(\mathbb{T}_{p})=\sum_{l^{\prime}=1}^{l}\big|S_{L(i,j)}^{(n)}(t_{r_{l^{\prime}}})-S_{L(i,j)}^{(n)}(t_{r_{l^{\prime}}-1})\big|$, $1\leq i,j\leq n$. Then,
$|S_{\pi}^{(n)}(\mathbb{T}_{p})|\leq\mathbb{E}M_{L(\pi(0),\pi(1))}^{(n)}(\mathbb{T}_{p})\dots M_{L(\pi(l-1),\pi(l))}^{(n)}(\mathbb{T}_{p}).$
For $h\geq 1$ and $s<t$, we have $\mathbb{E}|S_{L(i,j)}^{(n)}(t)-S_{L(i,j)}^{(n)}(s)|^{h}\leq K_{h}([nt]-[ns])^{[h/2]}$, which implies $\mathbb{E}(M_{L(i,j)}^{(n)}(\mathbb{T}_{p}))^{h}\leq\prod_{l^{\prime}=1}^{l}K_{h_{l^{\prime}}}([nt_{r_{l^{\prime}}}]-[nt_{r_{l^{\prime}}-1}])^{[h_{l^{\prime}}/2]}$, where $h_{1}+\dots+h_{l}=h$. Thus, for any non pair matched circuit $\pi$ of length $l$, following the arguments of the proof for Theorem 3.1, we get $|S_{\pi}^{(n)}(\mathbb{T}_{p})|\leq K_{l}\prod_{l^{\prime\prime}=1}^{l}([nt_{r_{l^{\prime}}}]-[nt_{r_{l^{\prime}}-1}])^{[h_{l^{\prime\prime}}^{\prime}/2]}$, where $h_{1}^{\prime}+\dots+h_{l}^{\prime}=l$. Now, on using $\Pi(w)\leq K_{l}n^{(l+1)/2}$ for any non pair-matched word $w$ of length $l$ (see [9], p. 120), we conclude that the contribution of a non indexed pair-matched circuit in (3.10) is zero as $n\rightarrow\infty$. Hence,
@@
where $\pi_{g,w}\in\Pi(\chi_{g}(w))$ is a fixed indexed pair-matched circuit of length $2m$. Moreover, we have
@@
Thus, from (3.5), we get $\tilde{\tau}((A(t)-A(s))^{l})=\tilde{\tau}((A(t-s))^{l})$ for all $l\geq 1$. If $L(i,j)=(\max\{i,j\},\min\{i,j\})$ then from (i) of Corollary 3.1, it follows that for all $s<t$, A(t)A(s)A(t)-A(s) is a semi-circle element with mean zero and variance tst-s. Moreover, in view of part (i), the variables in $\mathbb{T}_{p}(A(t))$, defined in (3.6), are freely independent using Theorem 9.2.1 of [9].
For the Symmetric circulant, similar arguments work. Finite dimensional convergence to moments of an SBM follows from Corollary 3.1 (v) and Theorem 9.5.1 of [9]. This completes the proof. ∎
@@
In [8], a general condition is obtained on the link function of random matrices that are not necessarily Wigner matrices, under which the associated LSD has the semi-circular density. Let $A_{n}(t)$ be the random matrix with link function $L$, as defined in Theorem 3.1. Define
$\beta_{n}=\max_{r}\#\{(i,j):L(i,j)=r,\ 1\leq i,j\leq n\},\ n\geq 1.$
@@
$M_{i,j}^{(n)}=\{k:1\leq k\leq n,L(k,i)=L(k,j)\},$
If $M^{*}=\sup_{n\geq 1}\sup_{1\leq i,j\leq n}\#M_{i,j}^{(n)}<\infty$, and $\beta_{n}=o(n)$, then in view of Theorem 2 of [8], the LSD of $A_{n}(t)$ has semi-circle density with variance $t>0$.
3.3 Symmetric circulant matrix and SBM
Now we consider $A_{n}(t)$’s with the link function L(i,j)=n/2|n/2|ij||L(i,j)=n/2-|n/2-|i-j||, $1\leq i,j\leq n$, and suppose Assumption I holds. Then $F_{n^{-1}A_{n}(t)},t\geq 0$, the ESD of $A_{n}(t),t\geq 0$ is a random process. We already know that for each $t>0$, the ESD $F_{n^{-1}A_{n}(t)}$ converges weakly a.s. to the Gaussian law with mean zero and variance $t$.
@@ 3.3 Symmetric circulant matrix and SBM
Let $\{W(t)\}_{t\geq 0}$ be an SBM, and let $\{Y_{n}(t)\}_{t\geq 0}$ be the random process, as defined in (3.11). Then, $\{Y_{n}(t)\}_{t\in[0,1]}$ converges weakly to $\{W(t)\}_{t\in[0,1]}$, as $n\rightarrow\infty$ in the Skorokhod space $D[0,1]$.
@@
4 Time-changed free Brownian motion
A Brownian motion time-changed by an independent inverse stable subordinator is a key mathematical model for analyzing anomalous diffusion in statistical mechanics. It is used for the modeling of motion of a particle experiencing a trapping effect. Here, we define a random time-changed FBM driven by an inverse stable subordinator. Let $\{B(t)\}_{t\geq 0}$ be an FBM on the *-probability space $(\mathcal{A},\tau)$. Let $L^{\alpha}(t),t\geq 0$ be an inverse stable subordinator on $(\Omega,\mathcal{F},\mathbb{P})$. From [31], all moments of $L^{\alpha}(t),t\geq 0$ are finite, Hence $\mathcal{A}_{\infty}$, the set of all polynomials in $L^{\alpha}(t),t\geq 0$ is an NCP with its state as expectation operator $\mathbb{E}$. Let us consider a collection $\{B^{\alpha}(t)\}_{t\geq 0}$ of variables of the tensor product space $(\hat{\mathcal{A}}:=\mathcal{A}\otimes\mathcal{A}_{\infty},\hat{\tau}:=\mathbb{E}\tau)$ defined as follows:
$B^{\alpha}(t)\coloneqq$ $B(L^{\alpha}(t)),\ 0<\alpha\leq 1,$ (4.1)
$\hat{\tau}(B^{\alpha}(t_{1})\dots B^{\alpha}(t_{l}))\coloneqq$ $\mathbb{E}\tau(B^{\alpha}(t_{1})\dots B^{\alpha}(t_{l})).$
$\hat{\tau}(B^{\alpha}(t_{1})\dots B^{\alpha}(t_{l})):=$ $\int_{\mathbb{R}^{l}_{+}}\tau(B(x_{1})\dots B(x_{l}))\mathrm{Pr}\{L^{\alpha}(t_{1})\in\mathrm{d}x_{1},\dots,L^{\alpha}(t_{l})\in\mathrm{d}x_{l}\}$
for all $l\geq 1$ and $t_{1},\dots,t_{l}\geq 0$. The collection $\{B^{\alpha}(t)\}_{t\geq 0}$ will be called a time-changed FBM. For $\alpha=1$, it reduces to the FBM.
where $C_{m}=(2m)!/(m+1)!m!$ is the $m$th Catalan number.
Proposition 4.1.
Let $Z(t)$ be a semi-circle rv with mean zero and variance $t\geq 0$ that is independent of an inverse $\alpha$-stable subordinator $\{L^{\alpha}(t)\}_{t\geq 0}$. Then,
$\mathbb{E}e^{uZ(L^{\alpha}(t))}=(u^{2}t^{\alpha})^{-1}\bigg(W_{\alpha,1-\alpha}(u^{2}t^{\alpha})-\frac{1}{\Gamma(1-\alpha)}\bigg),\ u\in\mathbb{R},\ t>0,$ (4.3)
where $W_{\nu,\mu}$, $\nu>-1$ is the Wright function (for definition see [18], Eq. (1.11.1)). As a consequence, $\{B^{\alpha}(t)\}_{t\geq 0}$ is a free process with $B^{\alpha}(0)=0$ and for each $t>0$, the spectral distribution of $B^{\alpha}(t)$ has a valid probability density with support $\mathbb{R}$, whose associated moment generating function (mgf) is given by (4.3).
Proof.
4.1 Patterned random matrices with randomly stopped walks
@@
Let $X_{1},X_{2},\dots$ be iid rvs that are independent of the fPP $\{N^{\alpha}(t)\}_{t\geq 0}$. Consider the scaled continuous time random walk $\{c^{-\alpha/2}\sum_{k=1}^{N^{\alpha}(ct)}X_{k}\}_{t\geq 0}$, $c>0$. It is known that if $X_{k}$’s belong to the domain of attraction of standard Gaussian distribution then $\{c^{-\alpha/2}\sum_{k=1}^{N^{\alpha}(ct)}X_{k}\}_{t\geq 0}$ weakly converges to a time-changed process $\{W(L^{\alpha}(t))\}_{t\geq 0}$ in an appropriate Skorokhod topology, as $c\rightarrow\infty$ (see [23]) where $\{W(t)\}_{t\geq 0}$ is an SBM that is independent of an inverse $\alpha$-stable subordinator $\{L^{\alpha}(t)\}_{t\geq 0}$ as in Section 2.4.
@@
In this section, we study the convergence of random matrices whose entries are randomly stopped independent random walks. Let $\{X_{i,j,k},\ i\geq 1,\ j\geq 1,\ k\geq 1\}$ be independent rvs that are independent of $\{N^{\alpha}(t)\}_{t\geq 0}$. We consider a bi-sequence $\{\mathcal{S}_{i,j,n}^{\alpha}(t),\ i\geq 1,\ j\geq 1\}_{t\geq 0}$, $n\in\mathbb{N}$ of continuous time random walks defined as follows:
$\mathcal{S}_{i,j,n}^{\alpha}(t)\coloneqq\sum_{k=1}^{N^{\alpha}(nt)}X_{i,j,k},\ t\geq 0,\ i\geq 1,\ j\geq 1,\ n\geq 1.$ (4.4)
@@
Note that (4.4) is a fractional compound Poisson process, and is a compound Poisson process for $\alpha=1$. Let us consider the patterned random matrices $\{A_{n}^{\alpha}(t)\}_{t\geq 0}$ defined as
$A_{n}^{\alpha}(t)\coloneqq n^{-\alpha/2}(\mathcal{S}^{\alpha}_{L(i,j),n}(t))_{1\leq i,j\leq n},\ 0<\alpha\leq 1,\ n\geq 1.$ (4.5)
@@
We also assume that $\{X_{i,j,k},\ 1\leq i\leq j,\ k\geq 1\}$ satisfies Assumption I and the link function $L$ of $A_{n}^{\alpha}(t)$, $0<\alpha\leq 1$ is symmetric, and satisfies Property B. Here also if $L$ takes its values in $\mathbb{N}$ then with abuse of notation the input sequence $\{X_{i,j,k},\ 1\leq i\leq j,\ k\geq 1\}$ is a bi-sequence $\{X_{i,k},\ i\geq 1,\ k\geq 1\}$.
@@
In commutative probability, a Markov process, time-changed by an inverse stable subordinator, inherently produces a heavy-tailed distribution. For example, the fPP, see Section 2.4. Expecting something similar, Figures 1 and 2 illustrate the ESDs of $n^{-1/2}A_{n}^{\alpha}(t)$ for different values of $\alpha\in(0,1]$, corresponding to Wigner and Symmetric Circulant type link functions, respectively. It is observed that, for smaller values of $\alpha$, the eigenvalue spectrum is more widely spread, whereas increasing $\alpha$ leads to a contraction in the range of the $x$-axis. That is, the tails of the ESDs diminish as $\alpha$ increases.
@@
For $\alpha=1$, entries of $A^{\alpha}_{n}(t)|_{\alpha=1}$ are $n^{-1/2}S_{L(i,j),n}(t)=n^{-1/2}\sum_{k=1}^{N(nt)}X_{L(i,j),k}$, where $\{N(t)\}_{t\geq 0}$ is a Poisson process. For each $1\leq i,j\leq n$, the increments in $\mathbb{T}_{p}(S_{L(i,j),n}(t))$ involve $X_{i,j,k}$’s indexed by Poisson points in disjoint intervals. Hence increments in $\mathbb{T}_{p}(S_{L(i,j),n}(t))$, as well as in $\mathbb{T}_{p}(A_{n}^{1}(t))$ are mutually independent. The following result on $\mathbb{T}_{p}(n^{-1/2}A_{n}^{1}(t))$ is similar to Theorem 3.3:
@@
$\mathbb{E}|S_{L(i,j),n}(t)-S_{L(i,j),n}(s)|^{l}$ $=\sum_{h\leq r=0}^{\infty}\mathbb{E}\Big|\sum_{k=h}^{r}X_{L(i,j),k}\Big|^{l}\mathbb{P}\{N(ns)=h,\,N(nt)=r\}$
@@
for every $l\geq 3$ and for all $1\leq i,j\leq n$, and $\mathbb{E}(S_{L(i,j),n}(t)-S_{L(i,j),n}(s))^{2}=n(t-s)$. Here, we have used the $l$th moment of Poisson process, that is, $\mathbb{E}(N(t))^{l}=\sum_{r=0}^{l}S(l,r)t^{r}$, to get the last step, and where $S(l,r)$ is the Stirling number of second kind. Now, the proof follows similar lines to that of Theorem 3.3 (i). ∎
@@
For $\alpha=1$, $S_{i,j,n}(t)=\sum_{k=1}^{N(nt)}X_{i,j,k}$, where $\{N(t)\}_{t\geq 0}$ is a homogeneous Poisson process independent of $\{X_{i,j,k}\}$, and $A^{1}_{n}(t)=n^{-1/2}(S_{L(i,j),n}(t))_{1\leq i,j\leq n}$. From Theorem 4.1, for $p=1$ and $t_{0}=0$, it follows by arguments similar to those in the proof of Theorem 3.2, that the ESD of $n^{-1/2}A^{1}_{n}(t)$ converges weakly a.s. to a probability distribution whose $l$th moment is given by
@@
If $L(i,j)=(\max\{i,j\},\min\{i,j\})$, then the limiting variables $\{A^{1}(t)\}_{t\geq 0}$ form an FBM. If L(i,j)=n/2|n/2|ij||L(i,j)=n/2-|n/2-|i-j||, $1\leq i,j\leq n$, then the finite dimensional moments of the limit variables agree with those of an SBM.
@@
Let $N(n)$ be a Poisson rv with mean $n$. Let $\{A_{k,n}\}_{k\geq 1}$ be a sequence of $n\times n$ independent random matrices defined as $A_{k,n}=n^{-1/2}(X_{L(i,j),k})_{1\leq i,j\leq n}$ for all $k\geq 1$, where the input sequence $X_{i,j,k}$’s are as in Theorem 3.1. Also, let $\{A_{k,n}\}_{k\geq 1}$ be independent of $N(n)$, and have a common symmetric link function, satisfying Property B. If the limits (3.4) exist, then the LSD of $n^{-1/2}(A_{1,n}+\dots+A_{N(n),n})$ coincides with that of $A_{k,n}$ for each $k\geq 1$. It has a probability law whose $l$th moment equals (4.7). Thus, from Remark 3.1, it follows that $n^{-1/2}(A_{1,n}+\dots+A_{N(n),n})$ and $n^{-1/2}(A_{1,n}+\dots+A_{n,n})$ have a common LSD.
Let $\{X_{i,j,k},1\leq i\leq j,\,k\geq 1\}$ be as in Theorem 3.1. Let us consider $S_{i,j,n}(t)=n^{-1/2}\sum_{k=1}^{N(nt)}X_{i,j,k}$ for $1\leq i,j\leq n$ and $t\geq 0$, and a random matrix $A_{n}(t)=(S_{L(i,j),n}(t))$. Then, it can be easily established that Theorem 4.1 holds true for the matrix valued process $\{A_{n^{\alpha}}(t)\}_{t\geq 0}$, $\alpha\in(0,1]$ for an appropriate link function $L$. We now use this fact and the next result to establish the finite dimensional algebraic convergence of $\{A^{\alpha}_{n}(t)\}_{t\geq 0}$ to the time-changed FBM $\{B^{\alpha}(t)\}_{t\geq 0}$ as defined in (4.1).

### v2

@@ 3.2 Joint algebraic convergence
$A_{n}(t)-A_{n}(s)\coloneqq\big((S_{L^{d}(i,j)}^{(n)}(t)-S_{L^{d}(i,j)}^{(n)}(s))\big)_{1\leq i,j\leq n},\ n\geq 1.$
As $\{S_{i,j}^{(n)}(t)\}_{t\geq 0}$’s have independent increments, $\{A_{n}(t)\}_{t\geq 0}$ also has independent increments.
@@
(ii) If $L^{2}(i,j)=(\max\{i,j\},\min\{i,j\})$ then (3.8) holds and the collection $\{A(t)\}_{t\geq 0}$ in $\tilde{\mathcal{A}}$ is a FBM. Similarly, if $L^{1}(i,j)=n/2-|n/2-|i-j||$, $1\leq i,j\leq n$ then the joint moments of $\{A(t)\}_{t\geq 0}$ agree with those of a BM.
@@
$S_{\pi}^{(n)}(\mathbb{T}_{p})=\mathbb{E}(S_{L^{d}(\pi(0),\pi(1))}^{(n)}(t_{r_{1}})-S_{L^{d}(\pi(0),\pi(1))}^{(n)}(t_{r_{1}-1}))\dots(S_{L^{d}(\pi(l-1),\pi(l))}^{(n)}(t_{r_{l}})-S_{L^{d}(\pi(l-1),\pi(l))}^{(n)}(t_{r_{l}-1})).$
@@
Note that if any index $r_{i}$, $i=1,\dots,l$ appears only once or circuit $\pi$ is non matched then by independent increments property and zero mean of $\{S_{i,j}^{(n)}(t)\}_{t\geq 0}$, we get $S_{\pi}^{(n)}(\mathbb{T}_{p})=0$. Thus, $S_{\pi}^{(n)}(\mathbb{T}_{p})=0$ for all non indexed matched circuits $\pi$. Let us write $M_{L^{d}(i,j)}^{(n)}(\mathbb{T}_{p})=\sum_{l^{\prime}=1}^{l}\big|S_{L^{d}(i,j)}^{(n)}(t_{r_{l^{\prime}}})-S_{L^{d}(i,j)}^{(n)}(t_{r_{l^{\prime}}-1})\big|$, $1\leq i,j\leq n$. Then,
$|S_{\pi}^{(n)}(\mathbb{T}_{p})|\leq\mathbb{E}M_{L^{d}(\pi(0),\pi(1))}^{(n)}(\mathbb{T}_{p})\dots M_{L^{d}(\pi(l-1),\pi(l))}^{(n)}(\mathbb{T}_{p}).$
For $h\geq 1$ and $s<t$, we have $\mathbb{E}|S_{L^{d}(i,j)}^{(n)}(t)-S_{L^{d}(i,j)}^{(n)}(s)|^{h}\leq K_{h}([nt]-[ns])^{[h/2]}$, which implies $\mathbb{E}(M_{L^{d}(i,j)}^{(n)}(\mathbb{T}_{p}))^{h}\leq\prod_{l^{\prime}=1}^{l}K_{h_{l^{\prime}}}([nt_{r_{l^{\prime}}}]-[nt_{r_{l^{\prime}}-1}])^{[h_{l^{\prime}}/2]}$, where $h_{1}+\dots+h_{l}=h$. Thus, for any non pair matched circuit $\pi$ of length $l$, following the arguments of the proof for Theorem 3.1, we get $|S_{\pi}^{(n)}(\mathbb{T}_{p})|\leq K_{l}\prod_{l^{\prime\prime}=1}^{l}([nt_{r_{l^{\prime}}}]-[nt_{r_{l^{\prime}}-1}])^{[h_{l^{\prime\prime}}^{\prime}/2]}$, where $h_{1}^{\prime}+\dots+h_{l}^{\prime}=l$. Now, on using $\Pi(w)\leq K_{l}n^{(l+1)/2}$ for any non pair-matched word $w$ of length $l$ (see [9], p. 120), we conclude that the contribution of a non indexed pair-matched circuit in (3.11) is zero as $n\rightarrow\infty$. Hence,
@@
where $\pi_{g,w}\in\Pi(\chi_{g}(w))$ is a fixed indexed pair-matched circuit of length $2m$. Moreover,
@@
Thus, from (3.6), we get $\tilde{\tau}((A(t)-A(s))^{l})=\tilde{\tau}((A(t-s))^{l})$ for all $l\geq 1$.
If $L^{2}(i,j)=(\max\{i,j\},\min\{i,j\})$ then from Corollary 3.1(i), it follows that for all $s<t$, A(t)A(s)A(t)-A(s) is a semi-circle element with mean zero and variance tst-s. Moreover, in view of part (i), the variables in $\mathbb{T}_{p}(A(t))$, defined in (3.7), are freely independent using Theorem 9.2.1 of [9].
For the Symmetric circulant, similar arguments work. Convergence of the moments of any monomial to the corresponding moments of a BM follows from Corollary 3.1 (v) and Theorem 9.5.1 of [9]. This completes the proof. ∎
@@
(i) In [8], a general condition is obtained on the link function of random matrices that are not necessarily Wigner matrices, under which the associated LSD has the semi-circular density. Let $A_{n}(t)$ be the random matrix with link function $L^{d}$, as defined in Theorem 3.1. Define
$\beta_{n}=\max_{r}\#\{(i,j):L^{d}(i,j)=r,\ 1\leq i,j\leq n\},\ n\geq 1.$
@@
$M_{i,j}^{(n)}=\{k:1\leq k\leq n,L^{d}(k,i)=L^{d}(k,j)\},$
If $M^{*}=\sup_{n\geq 1}\sup_{1\leq i,j\leq n}\#M_{i,j}^{(n)}<\infty$, and $\beta_{n}=o(n)$, then in view of Theorem 2 of [8], the LSD of $A_{n}(t)$ has semi-circle density with variance $t>0$. Indeed, Part (ii) of Theorem 3.3 for the corresponding process $\{A_{n}(t)\}_{t\geq 0}$ also holds in this case. We omit the proof.
(ii) Part (ii) of Theorem 3.3 can be extended in another direction. Consider $p$ independent copies of $\{A_{n}(t)\}_{t\geq 0}$ for the Wigner link function, say $L^{2}$ and the Symmetric circulant link function, say $L^{1}$. Then for $L^{2}$, these $p$ copies converge in *-distribution and are asymtotically free. For $L^{1}$, these $p$ copies converge in *-distribution and are asymptotically independent.
3.3 Symmetric circulant and weak convergence to a BM
Now we consider $A_{n}(t)$’s with the link function $L^{1}(i,j)=n/2-|n/2-|i-j||$, $1\leq i,j\leq n$, in more details. Then $F_{n^{-1}A_{n}(t)},t\geq 0$, the ESD of $A_{n}(t),t\geq 0$ is a random process. Suppose Assumption I holds. Then we already know that for each $t>0$, the ESD $F_{n^{-1}A_{n}(t)}$ converges weakly a.s. to the Gaussian law with mean zero and variance $t$.
@@ 3.3 Symmetric circulant and weak convergence to a BM
Let $\{W(t)\}_{t\geq 0}$ be a BM, and let $\{Y_{n}(t)\}_{t\geq 0}$ be the random process, as defined in (3.12). Then, $\{Y_{n}(t)\}_{t\geq 0}$ converges weakly to $\{W(t)\}_{t\geq 0}$, as $n\rightarrow\infty$.
@@
Remark 3.3.
As in Remark 3.2 (ii), Theorem 3.4 can be extended to $p$ independent copies of symmetric circulant matrices that satisfy Assumption I. Define the joint empirical process in the natural way by using the same $U_{n}$ for all the $p$ matrices. Then this process converges weakly a.s. to $p$ independent standard Brownnian motions. Later we shall study the usual circulant for which we prove similar results. Since those arguments are more general than what is required for the present case, we skip the details here.
4 Patterned matrices with randomly stopped walks
@@
Let $X_{1},X_{2},\dots$ be iid rvs that are independent of the fPP $\{N^{\alpha}(t)\}_{t\geq 0}$. Consider the scaled continuous time random walk $\{c^{-\alpha/2}\sum_{k=1}^{N^{\alpha}(ct)}X_{k}\}_{t\geq 0}$, $c>0$. It is known that if $X_{k}$’s belong to the domain of attraction of standard Gaussian distribution then $\{c^{-\alpha/2}\sum_{k=1}^{N^{\alpha}(ct)}X_{k}\}_{t\geq 0}$ weakly converges to a time-changed process $\{W(L^{\alpha}(t))\}_{t\geq 0}$ in an appropriate Skorokhod topology, as $c\rightarrow\infty$ (see [23]) where $\{W(t)\}_{t\geq 0}$ is a BM that is independent of an inverse $\alpha$-stable subordinator $\{L^{\alpha}(t)\}_{t\geq 0}$ as in Section 2.2.
@@
In this section, we study the convergence of random matrices whose entries are randomly stopped independent random walks. Let $\{X_{i,j,k},\ i\geq 1,\ j\geq 1,\ k\geq 1\}$ and $\{X_{i,k}\ i\geq 0,\ k\geq 1\}$ be collections of independent rvs that are independent of $\{N^{\alpha}(t)\}_{t\geq 0}$. We consider continuous time random walks defined as follows:
$\mathcal{S}_{i,j,n}^{\alpha}(t)\coloneqq$ $\sum_{k=1}^{N^{\alpha}(nt)}X_{i,j,k},\ t\geq 0,\ i\geq 1,\ j\geq 1,\ n\geq 1,$ (4.1)
$\mathcal{S}_{i,n}^{\alpha}(t)\coloneqq$ $\sum_{k=1}^{N^{\alpha}(nt)}X_{i,k},\ t\geq 0,\ i\geq 0,\ n\geq 1.$ (4.2)
@@
Note that (4.1) and (4.2) are fractional compound Poisson processes, and are compound Poisson processes for $\alpha=1$. Consider the patterned random matrices $\{A_{n}^{\alpha}(t)\}_{t\geq 0}$ as
$A_{n}^{\alpha}(t)\coloneqq n^{-\alpha/2}((\mathcal{S}^{\alpha}_{L^{d}(i,j),n}(t)))_{1\leq i,j\leq n},\ 0<\alpha\leq 1,\ n\geq 1.$ (4.3)
@@
We also assume that $\{X_{i,j,k},X_{i,k},\ 0\leq i\leq j,\ k\geq 1\}$ satisfies Assumption I and the link function $L^{d}$ is symmetric, and satisfies Property B.
@@
In commutative probability, a Markov process, time-changed by an inverse stable subordinator, inherently produces a process whose inter-event times have heavy tailed distributions. For example, the fPP, see Section 2.2. Expecting something similar, Figures 1 and 2 illustrate the ESDs of $n^{-1/2}A_{n}^{\alpha}(t)$ for different values of $\alpha\in(0,1]$, corresponding to Wigner and Symmetric Circulant type link functions, respectively. It is observed that, for smaller values of $\alpha$, the eigenvalue spectrum is more widely spread, whereas increasing $\alpha$ leads to a contraction in the range of the $x$-axis. That is, the tails of the ESDs diminish as $\alpha$ increases.
@@
For $\alpha=1$, entries of $A^{\alpha}_{n}(t)|_{\alpha=1}$ are $n^{-1/2}\mathcal{S}_{L^{d}(i,j),n}(t)=n^{-1/2}\sum_{k=1}^{N(nt)}X_{L^{d}(i,j),k}$, where $\{N(t)\}_{t\geq 0}$ is a Poisson process. For each $1\leq i,j\leq n$, the increments in $\mathbb{T}_{p}(\mathcal{S}_{L^{d}(i,j),n}(t))$ involve $X_{i,j,k}$’s or $X_{i,k}$’s indexed by Poisson points in disjoint intervals. Hence these increments as well as those in in $\mathbb{T}_{p}(A_{n}^{1}(t))$ are independent. The following result on $\mathbb{T}_{p}(n^{-1/2}A_{n}^{1}(t))$ is similar to Theorem 3.3:
@@
$\mathbb{E}|\mathcal{S}_{L^{d}(i,j),n}(t)-\mathcal{S}_{L^{d}(i,j),n}(s)|^{l}$ $=\sum_{h\leq r=0}^{\infty}\mathbb{E}\Big|\sum_{k=h}^{r}X_{L^{d}(i,j),k}\Big|^{l}\mathbb{P}\{N(ns)=h,\,N(nt)=r\}$
@@
for every $l\geq 3$ and for all $1\leq i,j\leq n$, and $\mathbb{E}(\mathcal{S}_{L^{d}(i,j),n}(t)-\mathcal{S}_{L^{d}(i,j),n}(s))^{2}=n(t-s)$. Here, we have used the $l$th moment of the Poisson process, that is, $\mathbb{E}(N(t))^{l}=\sum_{r=0}^{l}S(l,r)t^{r}$, to get the last step, and where $S(l,r)$ is the Stirling number of second kind. Now, the proof follows similar lines to that of Theorem 3.3 (i). ∎
@@
For $\alpha=1$, $\mathcal{S}^{1}_{i,j,n}(t)=\sum_{k=1}^{N(nt)}X_{i,j,k}$ and $\mathcal{S}^{1}_{i,n}(t)=\sum_{k=1}^{N(nt)}X_{i,k}$, where $\{N(t)\}_{t\geq 0}$ is a homogeneous Poisson process independent of $\{X_{i,j,k}\}$ and $\{X_{i,k}\}$, and $A^{1}_{n}(t)=n^{-1/2}(\mathcal{S}^{1}_{L^{d}(i,j),n}(t))_{1\leq i,j\leq n}$. From Theorem 4.1, for $p=1$ and $t_{0}=0$, it follows by arguments similar to those in the proof of Theorem 3.2, that the ESD of $n^{-1/2}A^{1}_{n}(t)$ converges weakly a.s. to a probability distribution whose $l$th moment is given by
@@
If $L^{2}(i,j)=(\max\{i,j\},\min\{i,j\})$, then the limiting variables $\{A^{1}(t)\}_{t\geq 0}$ form an FBM. If $L^{1}(i,j)=n/2-|n/2-|i-j||$, $1\leq i,j\leq n$, then the finite dimensional moments of the limit variables agree with those of a BM.
@@
Let $N(n)$ be a Poisson rv with mean $n$. Let $\{A_{k,n}\}_{k\geq 1}$ be a sequence of $n\times n$ independent random matrices defined as $A_{k,n}=n^{-1/2}((X_{L^{d}(i,j),k}))_{1\leq i,j\leq n}$ for all $k\geq 1$, where the input sequence $X_{i,j,k}$’s or $X_{i,k}$’s are as in Theorem 3.1. Also, let $\{A_{k,n}\}_{k\geq 1}$ be independent of $N(n)$, and have a common symmetric link function, satisfying Property B. If the limits (3.5) exist, then the LSD of $n^{-1/2}(A_{1,n}+\dots+A_{N(n),n})$ coincides with that of $A_{k,n}$ for each $k\geq 1$. It has a probability law whose $l$th moment equals (4.5). Thus, from Remark 3.1, it follows that $n^{-1/2}(A_{1,n}+\dots+A_{N(n),n})$ and $n^{-1/2}(A_{1,n}+\dots+A_{n,n})$ have a common LSD.
Let $\{X_{i,j,k},X_{i,k},\ i\geq 0,\ j,\,k\geq 1\}$ be as in Theorem 3.1. Let us consider
$S_{i,j,n}(t)=n^{-1/2}\sum_{k=1}^{N(nt)}X_{i,j,k},\ 1\leq i,j\leq n,\ t\geq 0,$
and a random matrix
$A_{n}(t)=((S_{L^{d}(i,j),n}(t)))_{1\leq i,j\leq n}.$
Then, it can be easily established that Theorem 4.1 holds true for the matrix valued process $\{A_{n^{\alpha}}(t)\}_{t\geq 0}$, $\alpha\in(0,1]$ for any appropriate link function $L^{d}$. We now use this fact and the next result to establish the algebraic convergence of $\{A^{\alpha}_{n}(t)\}_{t\geq 0}$ (see (4.3) for definition) to the time-changed FBM $\{B^{\alpha}(t)\}_{t\geq 0}$ as defined in (2.3).